OpenStudy (anonymous):

A force 10i-3j+6k Newton acts on a body of 5 kg and it displaces from 6i+5j-3kmetre to 10i-2j+7k metre.the work done is-

OpenStudy (agreene):

So, we know the starting displacement: <6,5,-3> and the ending displacement: <10,-2,7> We need to know how long the change in displacement is to calculate work. So, we need to use the distance formula. \[\sqrt{(i_2-i_1)^2+(j_2-j_1)^2+(k_2-k_1)^2}\] \[\sqrt{(10-6)^2+(-2-5)^2+(7--3)^2}\] \[\sqrt{4^2+(-7)^2+10^2}\] \[\sqrt{165}m\] Now we need to find the magnitude of the force: <10,-3,6> \[\sqrt{10^2+(-3)^2+6^2}=\sqrt{145}N\] Now we know the distance and the force applied, the eqn for work being: \[W=F*d\] \[W=\sqrt{145}\sqrt{165}=5\sqrt{957} J\approx154.6770J\]

OpenStudy (anonymous):

sorry agreena the options are a)100J b)0J c)121J d)no of this ans is 121 but I can't solved this but I also get the same ans